QUESTION 15 First Charter Bank Liabilities $2,000,000 Deposits $200,000 $1,800,000 $2,000,000 Assets Reserves Loans Total $2,QUESTION 16 $60,000 $40,000 Peoples Bank Assets Liabilities Total Reserves: $500,000 Deposits Required Owners Equity ExcessQUESTION 17 1 $150,000 $100,000 Liabilities $1,000,000 Deposits $200,000 Owners Equity Assets Total Reserves: Required Exces
Answer #1

Q15 Losses result in reduction in owners' equity, hence the right answer is the first one (even if the owners' equity becomes negative)

Q16 Since required reserve is 40,000 and deposit is 500,000, the required reserve ratio is:

40000 / 500000 = 8%

Q17 Total of assets = Total loans + total reserves

1200000 = x + 150000

x = 1200000 - 150000 = 1050000

Hence total loans = 1050000